# EndoGeDE: Generalizable Monocular Depth Estimation with Mixture of Low-Rank Experts for Diverse Endoscopic Scenes

### v2

@@ Abstract
Self-supervised monocular depth estimation is a significant task for low-cost and efficient 3D scene perception in endoscopy. In recent years, a series of methods are proposed to address the illumination inconsistency, while certain works also focus on the generalization of the model by efficiently finetuning the foundation models. However, the variety of illumination conditions and scene features is still the primary challenges for depth estimation in endoscopic scenes. In this work, a self-supervised framework is proposed for monocular depth estimation in diverse endoscopy. Firstly, considering the diverse features in endoscopic scenes with different tissues, a novel block-wise mixture of dynamic low-rank experts is proposed to efficiently finetune the foundation model for endoscopic depth estimation. In the proposed module, based on the input feature, different experts with a small amount of trainable parameters are adaptively selected for weighted inference, from low-rank experts which are allocated based on the generalization of each block. Moreover, a novel self-supervised training framework is proposed to jointly cope with brightness inconsistency and reflectance interference. The proposed method outperforms state-of-the-art works on SCARED dataset and SimCol dataset. Furthermore, the proposed network also achieves the best generalization based on zero-shot depth estimation on C3VD, Hamlyn and SERV-CT dataset. The outstanding performance of our model is further demonstrated with 3D reconstruction and ego-motion estimation. The proposed method could contribute to accurate endoscopy for minimally invasive measurement and surgery. The evaluation codes will be released upon acceptance, while the demo videos can be found on https://endo-gede.netlify.app/.
1 Introduction
@@
Endoscopy, as one of primary minimally invasive surgical procedures, is widely used in the medical treatment due to its minor additional trauma and short recovery period. However, due to the limited field of view and 2D vision of endoscopes, 3D scene perception has become a significant task for efficient and accurate endoscopy. Depth estimation, as one of the most common 3D vision techniques which has been applied in industrial scenes(Liu et al., 2024), underwater scenes(Ye et al., 2023), outdoor driving scenes(Feng et al., 2023; Zhu et al., 2025) and indoor scenes(Wei et al., 2024), has also been a primary 3D scene perception method in endoscopy.
@@
Due to the high cost of obtaining depth maps (Yang et al., 2024c) and limited inference efficiency of stereo matching (Wu et al., 2025) in endoscopy, self-supervised monocular depth estimation is the main-stream method for endoscopic scenes. Although numerous self-supervised methods, such as MonoViT(Zhao et al., 2022) and LiteMono(Zhang et al., 2023a), have been proposed for monocular depth estimation for natural scenes based on deep learning, there are two critical challenges for the application of such methods in endoscopy. Firstly, illumination conditions in endoscopic scenes often causes brightness and reflectance inconsistency. Moreover, the performance and generalization of the depth estimation model will be highly limited by the variety of scene features due to different surgical tasks and tissues, which is shown in Fig. 1.
@@
To deal with the illumination inconsistency, Shao et al.(Shao et al., 2022) propose to utilize constraint based on the appearance flow to alleviate brightness inconsistency. Meanwhile, intrinsic image decomposition is also a useful technique to filter the interference of the light, which is also applied in endoscopy(Li et al., 2024). However, as Fig. 1 shows, the depth at some positions with high brightness on the surface will be estimated as a higher value than the ground truth, while some light spots from reflectance on the surface will highly influence the depth value at the corresponding position. With the development of foundation models for depth estimation, a series of methods are proposed to efficiently finetune the foundation model for adaptation to the endoscopic scenes, including Surgical-DINO(Cui et al., 2024a), EndoDAC(Cui et al., 2024b) and DARES(Zeinoddin et al., 2024). However, most of the existing finetuned foundation models for endoscopic depth estimation are based on Low-Rank Adaptation (LoRA)(Hu et al., 2022), the performance of which may be constrained by the feature diversity in diverse endoscopic scenes.
To this end, a novel self-supervised framework, EndoGeDE, is proposed to boost the performance and generalization of monocular depth estimation in diverse endoscopic scenes. First of all, in the proposed framework, an end-to-end training pipeline based on the integration of a novel intrinsic image registration and brightness constraint (Shao et al., 2022) is utilized to jointly deal with brightness inconsistency and light interference. Specifically, the invariance of the reflectance image to illumination changes and the strong correlations in the gradient domain(Kim et al., 2016) are considered in the proposed intrinsic image decomposition. Furthermore, a novel parameter-efficient finetuning method, Block-wise Mixture of Low-Rank Experts (BW-MoLE), is proposed to finetune the foundation model for adaptive depth estimation in endoscopy based on diverse scene features, which is built upon LoRA and Mixture of Experts (MoE). Note that the low-rank experts are allocated based on the training quality of different Transformer blocks in the encoder of foundation model. The main contribution of this work can be summarized as the following:
@@
    The proposed method is evaluated by self-supervised depth estimation on SCARED(Allan et al., 2021) and SimCol(Rau et al., 2024) dataset, as well as zero-shot depth estimation on C3VD(Bobrow et al., 2023), Hamlyn(Recasens et al., 2021) and SERV-CT(Edwards et al., 2022) dataset. Our method achieves state-of-the-art performance compared with existing works. Moreover, the proposed method is applied to sim-to-real test(Azagra et al., 2023) and 3D reconstruction.
2 Related Work
@@
2.1 Self-supervised Depth Estimation for Endoscopy
@@
Based on the method proposed by Zhou et al.(Zhou et al., 2017), Turan et al. (Turan et al., 2018) proposed the first self-supervised method for depth estimation in endoscopic scenes. In the next few years, some works for self-supervised depth estimation in endoscopy are also proposed with the employment of attention mechanism. Ozyoruk et al. (Ozyoruk et al., 2021) introduced spatial attention blocks into the depth estimation network and ego-motion estimation network, while Liu et al. (Liu and Zuo, 2023) further integrated the spatial attention and channel attention in the proposed framework. In the most recent work, Yang et al. (Yang et al., 2024c) proposed a lightweight self-supervised depth estimation method for endoscopy with the combination of convolution and self-attention. Different from direct estimation from the observation, Liu et al. (Liu et al., 2023a), (Liu et al., 2023b) proposed a novel strategy to perform depth estimation based on a depth completion network, which reconstructs dense depth maps based on RGB data and corresponding sparse depth maps. Although a large amount of methods have been proposed for self-supervised depth estimation, the application of these methods in endoscopy is still limited by illumination issues in the endoscopic images.
@@
In 2022, Shao et al. (Shao et al., 2022) proposed AF-SfMLearner to firstly tackle the brightness inconsistency using a generalized dynamic image constraint based on the appearance flow in endoscopic scenes. Based on this, Li et al. (Li et al., 2024) recently proposed a self-supervised method based on intrinsic image decomposition to deal with the challenge of reflectance on the surface of the tissue for endoscopic depth estimation. Differently, Zhou et al. (Zhou et al., 2024) proposed a method to select candidate frames with lowest temporal variance and further assign the source frame with the highest view consistency. The most recent work, MonoPCC (Wang et al., 2025), utilized a novel photometric-invariant cycle constraint to obtain cycle-warped image from the target image to solve the problem of brightness inconsistency. In this work, a novel self-supervised training strategy is proposed to jointly deal with the illumination inconsistency and the reflectance interference in diverse endoscopic scenes.
2.2 Foundation Models for Endoscopic Depth Estimation
@@
With the development of the foundation models for depth estimation, such as DINO(Oquab et al., 2024), series of Depth Anything(Yang et al., 2024a),(Yang et al., 2024b), the application of such models in depth estimation for various scenes has gained increasing attention. For example, Yu et al.(Yu et al., 2024) utilize DINOv2 to integrate semantic and geometric features for depth estimation in outdoor scenes. Some methods for depth estimation in endoscopy based on foundation models are also proposed in recent years. Firstly, Cui et al. (Cui et al., 2024a) proposed to finetune DINO with Low-Rank Adaptation(LoRA) for depth estimation in endoscopic scenes. Based on this, a self-supervised framework, EndoDAC (Cui et al., 2024b), based on Depth Anything model (Yang et al., 2024a) with the novel parameter-efficient finetuning strategy named as dynamic-vector LoRA (DV-LoRA) and inserted convolution block is proposed to perform endoscopic depth estimation with generalizability. Meanwhile, Zeinoddin et al. (Zeinoddin et al., 2024) also proposed a different finetuning strategy named vector-LoRA which is utilized on the Depth Anything model. In this work, the foundation model for depth estimation in diverse endoscopic scenes is proposed with a novel parameter-efficient finetuning method based on the integration of MoE and LoRA, which can strengthen the adaptation of the model to different scene features.
@@
Figure 2: The pipeline of the proposed self-supervised learning.
3 Proposed Method
3.1 Overview
@@
The overview of the proposed method is shown as Fig. 2. The proposed framework consists of five networks, which respectively perform optical flow prediction, appearance flow prediction, depth map prediction, intrinsic image decomposition, ego-motion estimation and camera intrinsic estimation. Firstly, networks for optical flow prediction and appearance flow prediction are mainly from AF-SfMLearner(Shao et al., 2022), while ego-motion estimation and camera intrinsics estimation are based on EndoDAC(Cui et al., 2024b). The framework of depth map prediction is a finetuned encoder-decoder network based on Depth Anything(Yang et al., 2024a), which generates the corresponding depth map from single input image. Moreover, given an input observation, the reflectance image, the shading image and the soft assignment map will be generated by the proposed intrinsic image decomposition with an encoder-decoder framework.
@@
The self-supervised learning of the framework is divided into three stages in each end-to-end training epoch to reserve the computational resources and pretend the information interference, demonstrated in Section 3.4. The intrinsic image decomposition network is firstly trained based on the consistency between the recomposed images and the raw images at the second stage, which is introduced in Section 3.3. At the final stage, the depth map prediction network is efficiently finetuned by the proposed block-wise mixture of low-rank experts in Section 3.2. Meanwhile, other networks, except the optical flow prediction network trained at the first stage, are fully trained at the final stage based on the alignments of rigid transformations using predicted ego-motion and depth map. The photometric loss is utilized to supervise the consistency of images, defined as Eq. 1.
@@
where $V$ is the visibility mask generated from backward optical flow following (Shao et al., 2022), $\beta=0.85$.
3.2 Parameter-efficient Finetuning for Depth Map Prediction
@@
3.2.1 Mixture of Low-Rank Experts
Based on DV-LoRA(Cui et al., 2024b), multiple low-rank experts are implemented to update the output based on the frozen pretrained weights. Given the input feature, the route network $\mathcal{R}$ will predict the corresponding weights of the experts network by Eq. 2, in which the noise generated by the network $\mathcal{N}$ is also introduced for robustness. The experts with top $k$ weights $\alpha_{1},\alpha_{2},...\alpha_{k}\in\Gamma_{b}=\{\alpha_{1},\alpha_{2},...,\alpha_{E_{b}}|\alpha_{i}>\alpha_{i+1}\}$ are selected to perform weighted finetuning. The finetuning process can be formulated as Eq. 3.
@@ 3.2.1 Mixture of Low-Rank Experts
3.2.2 Block-wise Allocation of Experts
The weights of the network can be optimized by the smaller portion of specific features in the corresponding domain, which leads to limited generalization or transferring of the foundation model. Therefore, in this work, the number of experts is allocated based on the generalization of each block in the foundation model. To measure the generalizability, we analyze the distribution $\mathcal{D}(\lambda)$ of the eigenvalues $\lambda\in\Sigma^{2}$ based on the singular value decomposition $W=U\Sigma V$ of the extracted weights $W$ from linear layers, as Eq. 4 shows, where $\hat{\lambda}=argmax_{\lambda}(\log_{10}(\mathcal{D}(\lambda)))$. Based on the previous work(Zhou et al., 2023), the power law is utilized to fitting the probability distribution using the Hill estimator (Eq. 5) and the Fix-finger (Yang et al., 2023) method (Eq. 6).
@@ 3.2.2 Block-wise Allocation of Experts
Figure 4: The framework for intrinsic image decomposition
3.3 Self-supervised Loss for Intrinsic Image Decomposition
The framework for intrinsic image decomposition shown in Fig. 4 is designed based on the optical flow prediction and appearance flow prediction networks (Shao et al., 2022),(Li et al., 2024). Given the input image $I$, the albedo image $A$, the shading image $S$ and the assignment mask $M$ are predicted by the intrinsic image decomposition network. The albedo image represents the inherent color and texture of a surface independent of illumination conditions, while the shading image represents the light conditions. The assignment mask represents the probability of each pixel derivation from albedo, which can denote the influence of the illumination. The relationship of the raw image, the albedo image and the shading image can be defined as I=ASI=A*S.
Referring to (Kim et al., 2016), there are strong correlations between the raw image and the intrinsic images at the gradient domain. Based on Retinex algorithm(Land and McCann, 1971), the decomposed image can be estimated from the raw image and the assignment mask. To this end, the photometric loss between the gradient of the estimated decomposed image and the gradient of the predicted image will be utilized to supervise the optimization of the intrinsic image decomposition, formulated as Eq. 9.
3.4 Multi-stage Self-supervised Training
Each epoch of the end-to-end training is divided into three stages, including optical flow registration, self-supervised intrinsic image decomposition and multiple transformation alignments based on the depth map. In each step, the target image $I_{t}=I_{i}$ and the corresponding source images $I_{s}\in\{I_{i+n},I_{i-n}\}$ are given to the framework for optimization.
3.4.1 First Stage: Self-supervised Optical Flow Estimation
Following (Shao et al., 2022), the optical flow prediction network is optimized based on the loss function Eq. 10, where $O_{t}$ is the forward optical flow of the source image $I_{t}$. $I^{opt}_{s\rightarrow t}$ is generated from the source image $I_{t}$ based on the forward optical flow and Spatial Transformer.
3.4.2 Second Stage: Self-supervised Intrinsic Image Decomposition
At the second stage, the intrinsic image decomposition network is solely optimized by the loss function $\mathcal{L}_{2}$ shown as Eq. 11, consisting of the losses demonstrated in Section 3.3 based on the decomposition of the target image and the source images.
$\mathcal{L}_{2}=\mathcal{L}_{rec}+0.1\mathcal{L}_{ret}$ (11)
3.4.3 Third Stage: Multiple Transformation Alignments

### v3

# EndoGMDE: Generalizable Monocular Depth Estimation with Mixture of Low-Rank Experts for Diverse Endoscopic Scenes

@@ Abstract
Self-supervised monocular depth estimation is a significant task for low-cost and efficient 3D scene perception and measurement in endoscopy. However, the variety of illumination conditions and scene features is still the primary challenges for depth estimation in endoscopic scenes. In this work, a novel self-supervised framework is proposed for monocular depth estimation in diverse endoscopy. Firstly, considering the diverse features in endoscopic scenes with different tissues, a novel block-wise mixture of dynamic low-rank experts is proposed to efficiently finetune the foundation model for endoscopic depth estimation. In the proposed module, based on the input feature, different experts with a small amount of trainable parameters are adaptively selected for weighted inference, from low-rank experts which are allocated based on the generalization of each block. Moreover, a novel self-supervised training framework is proposed to jointly cope with brightness inconsistency and reflectance interference. The proposed method outperforms state-of-the-art works on SCARED dataset and SimCol dataset. Furthermore, the proposed network also achieves the best generalization based on zero-shot depth estimation on C3VD, Hamlyn and SERV-CT dataset. The outstanding performance of our model is further demonstrated with 3D reconstruction and ego-motion estimation. The proposed method could contribute to accurate endoscopy for minimally invasive measurement and surgery. The evaluation codes have been released on https://github.com/BaymaxShao/Endo-GMDE, while the demo videos can be found on https://endo-gmde.netlify.app/.
I Introduction
@@
Endoscopy, as one of primary minimally invasive surgical procedures, is widely used in the medical treatment due to its minor additional trauma and short recovery period. However, due to the limited field of view and 2D vision of endoscopes, 3D scene measurement has become a significant task for the precise perception of anatomical structures, such as lesion size and vascular boundaries.[1, 2] Depth estimation, as one of the most common 3D vision techniques which has been applied in underwater scenes[3, 4], outdoor driving scenes[5, 6] and indoor scenes[7], has also been a primary 3D scene perception method in minimally invasive surgery[2].
@@
Due to the high cost of obtaining depth maps [8] and limited inference efficiency of stereo matching [9] in endoscopy, self-supervised monocular depth estimation is the main-stream method for endoscopic scenes. Although numerous self-supervised methods, such as MonoViT[10] and LiteMono[11], have been proposed for monocular depth estimation for natural scenes based on deep learning, there are two critical challenges for the application of such methods in endoscopy. Firstly, illumination conditions in endoscopic scenes often causes brightness and reflectance inconsistency. Moreover, the performance and generalization of the depth estimation model will be highly limited by the variety of scene features due to different surgical tasks and tissues, which is shown in Fig. 1.
@@
To deal with the illumination inconsistency, Shao et al.[12] propose to utilize constraint based on the appearance flow to alleviate brightness inconsistency. Meanwhile, intrinsic image decomposition is also a useful technique to filter the interference of the light, which is also applied in endoscopy[13]. However, as Fig. 1 shows, the depth at some positions with high brightness on the surface will be estimated as a higher value than the ground truth, while some light spots from reflectance on the surface will highly influence the depth value at the corresponding position. With the development of foundation models for depth estimation, a series of methods are proposed to efficiently finetune the foundation model for adaptation to the endoscopic scenes, including Surgical-DINO[14], EndoDAC[15] and DARES[16]. However, most of the existing finetuned foundation models for endoscopic depth estimation are based on Low-Rank Adaptation (LoRA)[17], the performance of which may be constrained by the feature diversity in diverse endoscopic scenes.
To this end, a novel self-supervised framework, EndoGMDE, is proposed to boost the performance and generalization of monocular depth estimation in diverse endoscopic scenes. First of all, in the proposed framework, an end-to-end training pipeline based on the integration of a novel intrinsic image registration and brightness constraint [12] is utilized to jointly deal with brightness inconsistency and light interference. Specifically, the invariance of the reflectance image to illumination changes and the strong correlations in the gradient domain[18] are considered in the proposed intrinsic image decomposition. Furthermore, a novel parameter-efficient finetuning method, Block-wise Mixture of Low-Rank Experts (BW-MoLE), is proposed to finetune the foundation model for adaptive depth estimation in endoscopy based on diverse scene features, which is built upon LoRA and Mixture of Experts (MoE). Note that the low-rank experts are allocated based on the training quality of different Transformer blocks in the encoder of foundation model. The main contribution of this work can be summarized as the following:
@@
    The proposed method is evaluated by self-supervised depth estimation on SCARED[19] and SimCol[20] dataset, as well as zero-shot depth estimation on C3VD[21], Hamlyn[22] and SERV-CT[23] dataset. Our method achieves state-of-the-art performance compared with existing works. Moreover, the proposed method is applied to sim-to-real test[24] and 3D reconstruction.
II Related Work
@@
II-A Self-supervised Depth Estimation for Endoscopy
@@
Based on the method proposed by Zhou et al.[25], Turan et al. [26] proposed the first self-supervised method for depth estimation in endoscopic scenes. In the next few years, some works for self-supervised depth estimation in endoscopy are also proposed with the employment of attention mechanism. Ozyoruk et al. [27] introduced spatial attention blocks into the depth estimation network and ego-motion estimation network, while Liu et al. [28] further integrated the spatial attention and channel attention in the proposed framework. In the most recent work, Yang et al. [8] proposed a lightweight self-supervised depth estimation method for endoscopy with the combination of convolution and self-attention. Different from direct estimation from the observation, Liu et al. [29], [30] proposed a novel strategy to perform depth estimation based on a depth completion network, which reconstructs dense depth maps based on RGB data and corresponding sparse depth maps. Although a large amount of methods have been proposed for self-supervised depth estimation, the application of these methods in endoscopy is still limited by illumination issues in the endoscopic images.
@@
In 2022, Shao et al. [12] proposed AF-SfMLearner to firstly tackle the brightness inconsistency using a generalized dynamic image constraint based on the appearance flow in endoscopic scenes. Based on this, Li et al. [13] recently proposed a self-supervised method based on intrinsic image decomposition to deal with the challenge of reflectance on the surface of the tissue for endoscopic depth estimation. Differently, Zhou et al. [31] proposed a method to select candidate frames with lowest temporal variance and further assign the source frame with the highest view consistency. The most recent work, MonoPCC [32], utilized a novel photometric-invariant cycle constraint to obtain cycle-warped image from the target image to solve the problem of brightness inconsistency. In this work, a novel self-supervised training strategy is proposed to jointly deal with the illumination inconsistency and the reflectance interference in diverse endoscopic scenes.
II-B Foundation Models for Endoscopic Depth Estimation
@@
With the development of the foundation models for depth estimation, such as DINO[33], series of Depth Anything[34],[35], the application of such models in depth estimation for various scenes has gained increasing attention. For example, Yu et al.[36] utilize DINOv2 to integrate semantic and geometric features for depth estimation in outdoor scenes. Some methods for depth estimation in endoscopy based on foundation models are also proposed in recent years. Firstly, Cui et al. [14] proposed to finetune DINO with Low-Rank Adaptation(LoRA) for depth estimation in endoscopic scenes. Based on this, a self-supervised framework, EndoDAC [15], based on Depth Anything model [34] with the novel parameter-efficient finetuning strategy named as dynamic-vector LoRA (DV-LoRA) and inserted convolution block is proposed to perform endoscopic depth estimation with generalizability. Meanwhile, Zeinoddin et al. [16] also proposed a different finetuning strategy named vector-LoRA which is utilized on the Depth Anything model. In this work, the foundation model for depth estimation in diverse endoscopic scenes is proposed with a novel parameter-efficient finetuning method based on the integration of MoE and LoRA, which can strengthen the adaptation of the model to different scene features.
@@
Figure 2: The overview of this work and the pipeline of the proposed self-supervised training.
III Proposed Method
III-A Overview
@@
The overview of the proposed method is shown as Fig. 2. The proposed framework consists of five networks, which respectively perform optical flow prediction, appearance flow prediction, depth map prediction, intrinsic image decomposition, ego-motion estimation and camera intrinsic estimation. The framework of depth map prediction is a finetuned encoder-decoder network based on Depth Anything[34], which generates the corresponding depth map from single input image. Moreover, the ego-motion and camera intrinsics estimation network is mainly based on the module in [15].
@@
The self-supervised learning of the framework is divided into three stages in each end-to-end training epoch to reserve the computational resources and pretend the information interference. The optical flow estimation network is firstly trained based on self-supervised pipeline introduced in Section III-B. Then, the intrinsic image decomposition network is trained based on the consistency between the recomposed images and the raw images at the second stage, detail of which is introduced in Section III-C. At the final stage, the depth map prediction network is efficiently finetuned by the proposed block-wise mixture of low-rank experts in Section III-D. Meanwhile, other networks, except the optical flow estimation network trained at the first stage, are fully trained at the final stage based on the alignments of rigid transformations using estimated ego-motion and depth map, stated in Section III-E. The photometric loss is utilized to supervise the consistency of images, defined as Eq. 1.
@@
where $V$ is the visibility mask generated from backward optical flow following [12], $\beta=0.85$.
III-B Training Stage I: Optical Flow Registration
Following [12], the optical flow prediction network in Fig. 2 is optimized based on the loss function Eq. 2, where $O_{t}$ is the forward optical flow of the source image $I_{t}$. $I^{opt}_{s\rightarrow t}$ is generated from the source image $I_{t}$ based on the forward optical flow and Spatial Transformer.
III-C Training Stage II: Self-supervised Intrinsic Image Decomposition
Figure 3: The framework for intrinsic image decomposition
The framework for intrinsic image decomposition shown in Fig. 3 is designed based on the encoder-decoder structure for optical flow prediction and appearance flow prediction networks shown in Fig. 2. Given the input image $I$, the albedo image $A$, the shading image $S$ and the assignment mask $M$ are predicted by the intrinsic image decomposition network. The albedo image represents the inherent color and texture of a surface independent of illumination conditions, while the shading image represents the light conditions. The assignment mask represents the probability of each pixel derivation from albedo, which can denote the influence of the illumination. The relationship of the raw image, the albedo image and the shading image can be defined as I=ASI=A*S.
Referring to [18], there are strong correlations between the raw image and the intrinsic images at the gradient domain. Based on Retinex algorithm[37], the decomposed image can be estimated from the raw image and the assignment mask. To this end, the photometric loss between the gradient of the estimated decomposed image and the gradient of the predicted image will be utilized to supervise the optimization of the intrinsic image decomposition, formulated as Eq. 4.
At the second stage, the intrinsic image decomposition network is solely optimized by the loss function $\mathcal{L}_{2}$ shown as Eq. 5 based on above losses.
$\mathcal{L}_{2}=\mathcal{L}_{rec}+0.1\mathcal{L}_{ret}$ (5)
III-D Parameter-efficient Finetuning for Depth Map Prediction
@@
III-D1 Mixture of Low-Rank Experts
Based on DV-LoRA[15], multiple low-rank experts are implemented to update the output based on the frozen pretrained weights. Given the input feature, the route network $\mathcal{R}$ will predict the corresponding weights of the experts network by Eq. 6, in which the noise generated by the network $\mathcal{N}$ is also introduced for robustness. The experts with top $k$ weights $\alpha_{1},\alpha_{2},...\alpha_{k}\in\Gamma_{b}=\{\alpha_{1},\alpha_{2},...,\alpha_{E_{b}}|\alpha_{i}>\alpha_{i+1}\}$ are selected to perform weighted finetuning. The finetuning process can be formulated as Eq. 7.
@@ III-D1 Mixture of Low-Rank Experts
III-D2 Block-wise Allocation of Experts
The weights of the network can be optimized by the smaller portion of specific features in the corresponding domain, which leads to limited generalization or transferring of the foundation model. Therefore, in this work, the number of experts is allocated based on the generalization of each block in the foundation model. To measure the generalizability, we analyze the distribution $\mathcal{D}(\lambda)$ of the eigenvalues $\lambda\in\Sigma^{2}$ based on the singular value decomposition $W=U\Sigma V$ of the extracted weights $W$ from linear layers, as Eq. 8 shows, where $\hat{\lambda}=argmax_{\lambda}(\log_{10}(\mathcal{D}(\lambda)))$. Based on the previous work[38], the power law is utilized to fitting the probability distribution using the Hill estimator (Eq. 9) and the Fix-finger [39] method (Eq. 10).
@@ III-D2 Block-wise Allocation of Experts
III-E Stage III: Multiple Transformation Alignments